One worker make a wall in 5 days,second one in 6 days,3rd one in 7 days.if all three worker make wall together then how much time taken by them?
Answer by stanbon(75887)

One worker make a wall in 5 days,second one in 6 days,3rd one in 7 days.if all three worker make wall together then how much time taken by them?
-----
Rate Equation:
1/5 + 1/6 + 1/7 = 1/x
-----
Multiply thru by 210 to get:
42x + 35x + 30x = 210
----
107x = 210
---
x = 1.96 days (time to do the job together)
===========================
Cheers,
Stan H.